Summary of the Meeting of the PS Management Board
held on 05.11.2001
Present : B.W. Allardyce,
S. Baird, J. Boillot, M. Chanel, G. Daems, B. Frammery, R. Garoby, G. Geschonke,
J. Gruber, H. Haseroth, M. Lindroos, S. Maury, K. Metzmacher, S.L. Neboux (Secretary),
U. Raich, J.P. Riunaud, K. Schindl (Chairman)
Excused : D. Boimond (Staff Association), J.P. Delahaye, J.P. Potier
* * * * *
The meeting was chaired by K.
Schindl in the absence of J.P. Delahaye who was attending a special CERN Management
1. Follow-up of previous meeting and matters
- PSMB 29.10.2001 : no comments
- PS Days 2002 : due to the present circumstances at CERN, the
proposal to hold a second series of PS Days in 2002 in Evian or possibly
in Echenevex, has been cancelled.
- Safety :
- Road: due to the bad condition of the road surface and the
high expense of repair, G. Daems proposes to close the passage between
2008 and 175, to traffic circulation and to put concrete blocks
in place. The PSMB agreed to this proposal.
- EDH: the EDH request
for access authorisation, with automatic routing for signatures,
is now operational. This request is now linked to the attendance of
course. The access will be automatically cancelled if the course
is not followed in the following three weeks.
- CERN card and film : G. Daems has also proposed that the
wearing and display of the CERN
card and film
badge be made obligatory for people entering the PS tunnel.
This was also approved by the PSMB.
- Safety zone signs : R. Garoby was pleased to see that the
dilapidated signs had been replaced and thanked the DSO, G. Daems,
for this action.
- Replacing magnet units in the PS machine: J.P. Riunaud said that
in view of the forthcoming replacement of the magnet unit, the SL/MS
team are looking at the procedure for replacing units and to see if any
spare PS vacuum chambers are available.
- EPAC 2002 : the deadline
for abstract submission to EPAC is 11.01.2002. Therefore, the PS
Editorial Board needs to get them by 14 December 2001 at the
Associates Committee : the next meeting will take place on 13.11.2001.
Requests for 5 associates have been made and divisional priorities assigned,
but J. Boillot reminded the PSMB that the quota for associates is very
collaboration : S. Baird has distributed a memo to the PSMB concerning
a study to be carried out between the PS
Operation Group and the ST
Monitoring and Operation Group to improve collaboration between the
TCR and MCR after major breakdowns. It is an opportunity for the TCR to
know about our PS problems and needs, and vice-versa. A similar analysis
was made with SPS and found extremely useful by both sides. It will be
carried out in two phases : Linacs, Booster and PS during the 2001/2002
shutdown and ISOLDE, AD and CTF3 during the 2002/2003 shutdown. It will
essentially concern the Operation Group, but some information may also
be requested from other groups, especially from machine supervisors.
- ICALEPCS 2001
: B. Frammery announced that, for various reasons, only 2 members of CO
Group out of the 5 originally accepted will be attending this conference.
Due to recent events, only about one-third of participants will be attending
the conference, so a lot of the scientific interest is lost. It is proposed
to organise a CERN event to present papers prepared at CERN for ICALEPCS
- Schedule for end
of PS run : Schedule for end of PS run : S. Baird said that the
PS East Hall will stop at midnight today and the switch off of all PS
equipment will take place after the magnet tests, which will be carried
out on Tuesday morning. The PS radiation survey by TIS will take place
on Wednesday from 2-4 p.m. It is planned to run Linac2 and PSB for dedicated
MDs from Wednesday to Friday p.m. Interventions on power supplies, etc.
can take place on Friday as from 4 p.m. There will be no beam for ISOLDE
until Monday, 12.11.01. The PS ring will stay closed until the end of
the PSB/ISOLDE run on 3.12.2001.
- Reduced schedule meetings : meetings will continue on Mondays
for ISOLDE, MD results and CTF - PSMB members are welcome to attend.
- ISOLDE front-end: M. Lindroos reported that the new front-end
will be taken out this week in order to analyse the current leak problem.
Although it is not possible to be sure yet, oil contamination is suspected
to be the problem. Very likely no GPS operation will be possible this
year and only operation with HRS will continue.
: the new REX linear accelerator produced radioactive ions (Na26)
at 2 MeV/n for the first time. This device is unique because it is based
on a novel charge-breeding scheme. Even a first experiment was performed
over the weekend. K. Schindl, on behalf of PS division, congratulated
the whole team for the remarkable success of having accelerated radioactive
ions, after charge-breeding, for the first time. Besides ISOLDE-REX, two
other installations are also working at GANIL
(Spiral) and TRIUMF
(ISAC), and useful
collaborations are carried out between the three.
3. Report from the SPSC
J.P. Riunaud gave a summary
of the 55th SPSC meeting held on 30.10.2001. Items of particular interest
to PS concerned the status of SPS, PS and their experiments, reports from
HARP and DIRAC,
and the schedules proposed for 2002. The SPSC asked for the 25 ns run requested
for 2002 to be moved to 2003 and to start heavy ions a few days earlier
than foreseen (week 37). The proposed schedule foresees a reduction of PS
physics by 21 days and SPS physics by 7 days and gives rise to a saving
of some 4 MCHF. At first sight, the schedule appears to be acceptable to
the physics community, but any further cuts in beam allocation should be
avoided. The committee also expressed their concern about the reduced consolidation
of PS with hardly any reserve equipment.
4. Second job of operators during the shutdown
An analysis of second jobs for operators has been carried out in OP Group
and S. Baird gave a presentation
of first observations and ideas. During the 2000/2001 shutdown, operators
worked on software applications, hardware development and maintenance, training
of operators, pole-face winding measurements and safety procedures.
During the 2001/2002 shutdown, there will be 4 months without shift and
in 2002/2003 about 6 months. Apart from some time spent on specific training
on accelerator operation, resources will be available for other groups.
S. Baird presented the situation for the 2001/2002 shutdown and areas where
help would be available. Several jobs have to be continued and a few requests
for jobs have been received so far. Group and Project Leaders should contact
S. Baird to let him know their needs as soon as possible. For software jobs
B. Frammery requested that CO be informed of such requests since it will
have to provide some kind of support in any case. Action : Group and
He stressed also that for operators taking on a new second job,
it is vital that groups give clear specifications, as well as sufficient
technical help and supervision. M. Lindroos added that there are good examples
of excellent supervision and productivity, where second jobs were of benefit
to the operators and to the groups concerned. It was also proposed to think
ahead and arrange some brainstorming about second jobs during the longer
shutdowns and to try to improve the procedure.
| PS Division | PSMB
Meetings | Other PS Meetings |
6 November, 2001